OpenStudy (anonymous):

Diane needs 165 pieces of candy to hand out to guests at a party she is hosting. Each bag of candy contains 20 pieces. She has 25 pieces left over from a previous party. She calculates she needs to purchase 7 additional bags of candy to meet the 165 piece of candy requirement. Which of the equations below verifies that she needs 7 additional bags of candy?

OpenStudy (anonymous):

are there choices

OpenStudy (anonymous):

20.7+25=165

OpenStudy (anonymous):

Each bag has 20 pieces of candy, right?

OpenStudy (anonymous):

Since she figured she needs 7 bags, how many pieces of candy will she get in 7 bags? What operation would you use to find how many pieces of candy are in 7 bags of 20 pieces each?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

1 bag = 1 * 20 pieces 2 bags = 2 * 20 pieces 3 bags = 3 * 20 pieces ... 7 bags =

OpenStudy (anonymous):

I meant above for you to complete the last line. 7 bags =

OpenStudy (anonymous):

1 bag = 1 * 20 pieces 2 bags = 2 * 20 pieces 3 bags = 3 * 20 pieces ... 7 bags = 7 * 20 pieces

OpenStudy (anonymous):

The 7 bags she figured she needs to buy contain 7 * 20 pieces of candy. She already has 25 pieces of candy from before, so she needs to add the number of pieces of candy she gets in the 7 bags to the 25 she already had to see if that equals the 165 pieces she needs. 7 * 20 + 25 = 165

OpenStudy (anonymous):

That was my explanation. Just go over it and try to understand it. If you questions, ask, and I'll try to answer when I get back.

OpenStudy (anonymous):

@Its_James08

OpenStudy (anonymous):

can i get a medal